A wire having the diameter of 1.2 mm and resistively of 100 * 10 ^ - 1 * Omega * m at 0°C is connected. across a cell of emf


1.5 V and gives a current of 10 mA. Calculate the length of the wire.


1
Expert's answer
2022-03-06T15:19:57-0500

We know that

V=iR

"R=\frac{V}{I}"

"R=\frac{1.5}{0.01}=150\Omega"

We know that

"R=\frac{\rho l}{A}"

"l=\frac{RA}{\rho}"


"l=\frac{150\times\pi\times 36\times10^{-8}}{10}=1.69\times10^{-5}m"